Hi I'm having trouble in with my homework trying to find x, I'm really confused heres one of the problems
1/2(2x+54)= 1/3(9x+27)-4
A STEP BY STEP explanation would be greatly appreciated
Thanks, Ricky

Answer by stanbon(75887)

(1/2)(2x+54) = (1/3)(9x+27)-4
x+27 = 3+9-4
x+27 = 8
x = -19
==============
Cheers,
Stan H.
=========